# On total transitivity of graphs

### v1

@@ Abstract
Let $G=(V,E)$ be a graph where $V$ and $E$ are the vertex and edge sets, respectively. For two disjoint subsets $A$ and $B$ of $V$, we say $A$ dominates $B$ if every vertex of $B$ is adjacent to at least one vertex of $A$. A vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ of $G$ is called a transitive partition of size $k$ if $V_{i}$ dominates $V_{j}$ for all $1\leq i<j\leq k$. In this article, we study a variation of transitive partition, namely total transitive partition. The total transitivity $Tr_{t}(G)$ is equal to the maximum order of a vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ of $G$ obtained by repeatedly removing a total dominating set from $G$, until no vertices remain. Thus, $V_{1}$ is a total dominating set of $G$, $V_{2}$ is a total dominating set of the graph $G_{1}=G-V_{1}$, and, in general, $V_{i+1}$ is a total dominating set in the graph $G_{i}=G-\bigcup_{j=1}^{i}V_{i}$. A vertex partition of order $Tr_{t}(G)$ is called $Tr_{t}$-partition. The Maximum Total Transitivity Problem is to find a total transitive partition of a given graph with the maximum number of parts. First, we characterize split graphs with total transitivity equal to $1$ and $\omega(G)-1$. Moreover, for the split graph $G$ and $11\leq p\leq\omega(G)-$ 1, we give some necessary conditions for $Tr_{t}(G)=p$. Furthermore, we show that the decision version of this problem is NP-complete for bipartite graphs. On the positive side, we design a polynomial-time algorithm to solve Maximum Total Transitivity Problem in trees.
Keywords. Total transitivity, Split graphs, NP-completeness, Polynomial-time algorithm, Trees.
@@ 1 Introduction
Partitioning a graph is one of the fundamental problems in graph theory. In the partitioning problem, the objective is to partition the vertex set (or edge set) into some parts with desired properties, such as independence, minimal edges across partite sets, etc. In literature, partitioning the vertex set into certain parts so that the partite sets follow particular domination relations among themselves has been studied [1, 6, 8, 9, 13, 17, 18, 24]. Let $G$ be a graph with $V(G)$ as its vertex set and $E(G)$ as its edge set. When the context is clear, $V$ and $E$ are used instead of $V(G)$ and $E(G)$. The neighbourhood of a vertex $v\in V$ in a graph $G=(V,E)$ is the set of all adjacent vertices of $v$ and is denoted by $N_{G}(v)$. The degree of a vertex $v$ in $G$, denoted as $\deg_{G}(v)$, is the number of edges incident to $v$. A vertex $v$ is said to dominate itself and all its neighbouring vertices. A dominating set of $G=(V,E)$ is a subset of vertices $D$ such that every vertex $x\in V\setminus D$ has a neighbour $y\in D$, that is, $x$ is dominated by some vertex $y$ of $D$.
@@
There has been a lot of research on graph partitioning problems based on a domination relationship between the different sets. Cockayne and Hedetniemi introduced the concept of domatic partition of a graph $G=(V,E)$ in 1977, in which the vertex set is partitioned into $k$ parts, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, such that each $V_{i}$ is a dominating set of $G$ [4]. The number representing the highest possible order of a domatic partition is called the domatic number of G, denoted by $d(G)$. Another similar type of partitioning problem is Grundy partition. Christen and Selkow introduced a Grundy partition of a graph $G=(V,E)$ in 1979 [2]. In the Grundy partitioning problem, the vertex set is partitioned into $k$ parts, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, such that each $V_{i}$ is an independent set and for all $1\leq i<j\leq k$, $V_{i}$ dominates $V_{j}$. The maximum order of such a partition is called Grundy number of $G$, denoted by $\Gamma(G)$. In 2018, Hedetniemi and Hedetniemi [13] introduced a transitive partition as a generalization of the Grundy partition. A transitive partition of size $k$ is defined as a partition of the vertex set into $k$ parts, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, such that for all $1\leq i<j\leq k$, $V_{i}$ dominates $V_{j}$. The maximum order of such a transitive partition is called transitivity of $G$ and is denoted by $Tr(G)$. Recently, in 2020, Haynes et al. generalized the idea of domatic partition as well as transitive partition and introduced the concept of upper domatic partition of a graph $G$, where the set of the vertex is divided into $k$ parts, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, such that for each $i,j$, with $1\leq i<j\leq k$, either $V_{i}$ dominates $V_{j}$ or $V_{j}$ dominates $V_{i}$ or both [9]. The maximum order of such an upper domatic partition is called upper domatic number of $G$, denoted by $D(G)$. All these problems—domatic number [1, 26, 27], Grundy number [5, 6, 14, 24, 25], transitivity [8, 13, 17, 18], upper domatic number [9, 23] have been extensively studied both from an algorithmic and structural point of view. A Grundy partition is a transitive partition with the additional restriction that each partite set must be independent. In a transitive partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ of $G$, we require the domination property in one direction, that is, $V_{i}$ dominates $V_{j}$ for $1\leq i<j\leq k$. However, in an upper domatic partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ of $G$, for all $1\leq i<j\leq k$, either $V_{i}$ dominates $V_{j}$ or $V_{j}$ dominates $V_{i}$ or both. The definition of each vertex partitioning problem ensures the following inequalities for any graph $G$. For any graph $G$, $1\leq\Gamma(G)\leq Tr(G)\leq D(G)\leq n$, where $n$ is the number of vertices of $G$.
@@
Motivation: Many variations of the classical domination problem have been studied in the literature [10, 11, 12]. One of the natural variations is total dominating set, where our goal is to find a set $D$ such that every vertex of $V$ is dominated by a vertex of $D$. The concept of total domination was first formally introduced in 1980 by Cockayne et al. [3] as part of their study on domination in graphs. In the classical domination problem, we are interested in monitoring the vertices of a network that are not in a dominating set by the vertices of the dominating set. Total dominating ensures that every node in a network is adjacent to a control node, making it useful for designing fault-tolerant systems where every node has a backup connection. This variation of the domination problem has been well studied in the literature from both a structural and algorithmic point of view. It has been shown that finding the minimum total dominating set is NP-complete for general graphs, bipartite graphs [22], and split graphs [15], whereas it is solvable in polynomial time for trees [16].
@@
The domatic number of a graph was introduced by E. J. Cockayne and S. T. Hedetniemi [4]. Furthermore, many variants of this concept were introduced and studied by B. Zelinka [30]. One such variation is total domatic number, which was introduced by Cockayne et al. [3]. There are several works done on total domatic number [28, 29]. In 2018, Hedetniemi and Hedetniemi [13] introduced the notion of transitive partition. In 2024, S. Paul and K. Santra introduced three variations of transitivity, namely, $2$-transitivity [20], $d_{2}$-transitivity [19], and strong transitivity [21]. The $2$-transitivity, $d_{2}$-transitivity, and strong transitivity are based on $2$-domination, $d_{2}$-domination, and strong domination, respectively. In this article, we explore another natural variation of the transitivity problem, focusing on total domination instead of domination. We introduce such a variation of transitivity problem, namely total transitive partition. Total transitivity $Tr_{t}(G)$ equals the maximum order of a vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ of $G$ obtained by repeatedly removing a total dominating set from $G$, until no vertices remain. Thus, $V_{1}$ is a total dominating set of $G$, $V_{2}$ is a total dominating set of the graph $G_{1}=G-V_{1}$, and, in general, $V_{i+1}$ is a total dominating set in the graph $G_{i}=G-\bigcup_{j=1}^{i}V_{i}$. A vertex partition of order $Tr_{t}(G)$ is called $Tr_{t}$-partition. Note that every total transitive partition is also a transitive partition. Therefore, for any graph $G$, $1\leq Tr_{t}(G)\leq Tr(G)\leq n$. Since the total dominating set exists only when $G$ does not have any isolate vertex, in this paper, we focus only on graphs with no isolates.
@@
An equivalent definition of total transitivity can be given in terms of domination among partite sets and the set itself. For two disjoint subsets $A$ and $B$, we say $A$ dominates $B$ if every vertex of $B$ is adjacent to a vertex of $A$. Again, a subset $A$ dominates itself if every vertex of $A$ is adjacent to another vertex of $A$. A total transitive partition of size $k$ is defined as a partition of the vertex set into $k$ parts, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, such that for all $1\leq i\leq j\leq k$, $V_{i}$ dominates $V_{j}$. The maximum order of such a total transitive partition is called the total transitivity of $G$ and is denoted by $Tr_{t}(G)$.
@@
We define another variation of transitivity, namely modified-total transitivity. For two disjoint subsets $A$ and $B$, we say $A$ m-totally dominates $B$ if every vertex of $A\cup B$ is adjacent to a vertex of $A$. A modified-total transitive partition of size $k$ is defined as a partition of the vertex set into $k$ parts, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, such that for all $1\leq i<j\leq k$, $V_{i}$ m-totally dominates $V_{j}$. The maximum order of such a modified-total transitive partition is called the modified-total transitivity of $G$ and is denoted by $Tr_{t}^{m}(G)$. Later, we will use this parameter to solve total transitivity in a tree. Clearly, if $\pi$ is a total transitive partition, it is also a modified-total transitive partition, and hence, for a graph $G$, $Tr_{t}(G)\leq Tr_{t}^{m}(G)$. Also, note that in a modified-total transitive $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, $V_{k}$ can contain a single vertex, but if $\pi$ is a total transitive partition, $|V_{k}|\geq 2$.
@@
    Given a split graph $G$ and integer $p$, $11\leq p\leq\omega(G)-$ 1, we give some necessary conditions for $Tr_{t}(G)=p$.
@@
    We show that the Maximum Total Transitivity Problem can be solved in polynomial time for trees.
The rest of the paper is organized as follows. Section 2 contains basic definitions and notations that are followed throughout the article. This section also discusses the properties of total transitivity of graphs. In section 3, we characterize split graphs with total transitivity equal to $1$ and $\omega(G)-1$. Also, in this section, we give some necessary conditions for $Tr_{t}(G)=p$ for split graphs $G$ and $11\leq p\leq\omega(G)-$ 1. Section 4 shows that the Maximum Total Transitivity Decision Problem is NP-complete in bipartite graphs. In Section 5, we design a polynomial-time algorithm for solving Maximum Total Transitivity Problem in trees. Finally, Section 6 concludes the article with some open problems.
@@ 2.1 Notation and definition
Let $G=(V,E)$ be a graph with $V$ and $E$ as its vertex and edge sets, respectively. A graph $H=(V^{\prime},E^{\prime})$ is said to be a subgraph of a graph $G=(V,E)$ if and only if $V^{\prime}\subseteq V$ and $E^{\prime}\subseteq E$. For a subset $S\subseteq V$, the induced subgraph on $S$ of $G$ is defined as the subgraph of $G$ whose vertex set is, $S$ and edge set consists of all of the edges in $E$ that have both endpoints in $S$, and it is denoted by $G[S]$. The complement of a graph $G=(V,E)$ is the graph $\overline{G}=(\overline{V},\overline{E})$, such that $\overline{V}=V$ and $\overline{E}=\{uv|uv\notin E\text{ and }u\neq v\}$. The open neighbourhood of a vertex $x\in V$ is the set of vertices $y$ adjacent to $x$, denoted by $N_{G}(x)$. The closed neighborhood of a vertex $x\in V$, denoted as $N_{G}[x]$, is defined by $N_{G}[x]=N_{G}(x)\cup\{x\}$. A subset $D\subseteq V$ is said to be a total dominating set of $G$, if for every vertex $x\in V$, there exists a vertex $y\in D$ such that $xy\in E$. Breadth first search (BFS) is a graph searching technique that systematically explores all the vertices of a connected graph. The process starts at a designated vertex and explores all its neighbours before moving on to the next level of neighbours. The ordering in which the vertices are visited is called BFS ordering. A vertex is said to be a leaf vertex if its degree is $1$, and the vertex that is adjacent to a leaf vertex is said to be the support vertex of the leaf vertex.

### v2

@@ Abstract
Let $G=(V,E)$ be a graph where $V$ and $E$ are the vertex and edge sets, respectively. For two disjoint subsets $A$ and $B$ of $V$, we say that $A$ dominates $B$ if every vertex of $B$ is adjacent to at least one vertex of $A$. A vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ of $G$ is called a transitive partition of size $k$ if $V_{i}$ dominates $V_{j}$ for all $1\leq i<j\leq k$. In this article, we study a variation of the transitive partition, namely the total transitive partition. The total transitivity $Tr_{t}(G)$ is defined as the maximum order of a vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ of $G$ obtained by repeatedly removing a total dominating set from $G$ until no vertices remain. Thus, $V_{1}$ is a total dominating set of $G$, $V_{2}$ is a total dominating set of the graph $G_{1}=G-V_{1}$, and, for $12\leq i\leq k-$ 1, $V_{i+1}$ is a total dominating set in the graph $G_{i}=G-\bigcup_{j=1}^{i}V_{j}$. A vertex partition of order $Tr_{t}(G)$ is called a $Tr_{t}$-partition. The Maximum Total Transitivity Problem is to find a total transitive partition of a given graph with the maximum number of parts. First, we characterize split graphs with total transitivity equal to $1$ and $\omega(G)-1$. Moreover, for a split graph $G$ and $11\leq p\leq\omega(G)-$ 1, we provide necessary conditions for $Tr_{t}(G)=p$. Furthermore, we show that the decision version of this problem is NP-complete for bipartite graphs. On the positive side, we prove that this problem can be solved in linear time for bipartite chain graphs. Finally, we design a polynomial-time algorithm to solve the Maximum Total Transitivity Problem in trees.
Keywords. Total transitivity, Split graphs, NP-completeness, Bipartite chain graphs, Polynomial-time algorithm, Trees.
@@ 1 Introduction
Partitioning a graph is one of the fundamental problems in graph theory. In the partitioning problem, the objective is to partition the vertex set (or edge set) into some parts with desired properties, such as independence, minimal edges across partite sets, etc. In literature, partitioning the vertex set into certain parts so that the partite sets follow particular domination relations among themselves has been studied [1, 6, 8, 9, 13, 18, 19, 25]. Let $G$ be a graph with $V(G)$ as its vertex set and $E(G)$ as its edge set. When the context is clear, $V$ and $E$ are used instead of $V(G)$ and $E(G)$. The neighbourhood of a vertex $v\in V$ in a graph $G=(V,E)$ is the set of all adjacent vertices of $v$ and is denoted by $N_{G}(v)$. The degree of a vertex $v$ in $G$, denoted as $\deg_{G}(v)$, is the number of edges incident to $v$. A vertex $v$ is said to dominate itself and all its neighbours. A dominating set of $G=(V,E)$ is a subset of vertices $D$ such that every vertex $x\in V\setminus D$ has a neighbour $y\in D$, that is, $x$ is dominated by some vertex $y$ of $D$.
@@
There has been a lot of research on graph partitioning problems based on a domination relationship between the different sets. Cockayne and Hedetniemi introduced the concept of domatic partition of a graph $G=(V,E)$ in 1977, in which the vertex set is partitioned into $k$ parts, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, such that each $V_{i}$ is a dominating set of $G$ [4]. The number representing the highest possible order of a domatic partition is called the domatic number of G, denoted by $d(G)$. Another similar type of partitioning problem is Grundy partition. Christen and Selkow introduced a Grundy partition of a graph $G=(V,E)$ in 1979 [2]. In the Grundy partitioning problem, the vertex set is partitioned into $k$ parts, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, such that each $V_{i}$ is an independent set and for all $1\leq i<j\leq k$, $V_{i}$ dominates $V_{j}$. The maximum order of such a partition is called Grundy number of $G$, denoted by $\Gamma(G)$. In 2018, Hedetniemi and Hedetniemi [13] introduced a transitive partition as a generalization of the Grundy partition. A transitive partition of size $k$ is defined as a partition of the vertex set into $k$ parts, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, such that for all $1\leq i<j\leq k$, $V_{i}$ dominates $V_{j}$. The maximum order of such a transitive partition is called transitivity of $G$ and is denoted by $Tr(G)$. In 2020, Haynes et al. generalized the idea of domatic partition as well as transitive partition and introduced the concept of upper domatic partition of a graph $G$, where the set of the vertex is divided into $k$ parts, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, such that for each $i,j$, with $1\leq i<j\leq k$, either $V_{i}$ dominates $V_{j}$ or $V_{j}$ dominates $V_{i}$ or both [9]. The maximum order of such an upper domatic partition is called upper domatic number of $G$, denoted by $D(G)$. All these problems, domatic number [1, 27, 28], Grundy number [5, 6, 14, 25, 26], transitivity [8, 13, 18, 19], upper domatic number [9, 24] have been extensively studied both from an algorithmic and structural point of view. A Grundy partition is a transitive partition with the additional restriction that each partite set must be independent. In a transitive partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ of $G$, we require the domination property in one direction, that is, $V_{i}$ dominates $V_{j}$ for $1\leq i<j\leq k$. However, in an upper domatic partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ of $G$, for all $1\leq i<j\leq k$, either $V_{i}$ dominates $V_{j}$ or $V_{j}$ dominates $V_{i}$ or both. The definition of each vertex partitioning problem ensures the following inequalities for any graph $G$. For any graph $G$, $1\leq\Gamma(G)\leq Tr(G)\leq D(G)\leq n$, where $n$ is the number of vertices of $G$.
1.1 Motivation
@@
Many variations of the classical domination problem have been studied in the literature [10, 11, 12]. One natural variation is the total dominating set, where the goal is to find a set $D$ such that every vertex in $V$ is dominated by a vertex in $D$. The concept of total domination was first formally introduced in 1980 by Cockayne et al. [3] as part of their study on domination in graphs.In the classical domination problem, the objective is to monitor the vertices of a network that are not in a dominating set by using the vertices of the dominating set. Total domination ensures that every node in a network is adjacent to a control node, making it useful for designing fault-tolerant systems where every node has a backup connection.This variation of the domination problem has been well-studied in the literature from both structural and algorithmic perspectives. It has been shown that finding the minimum total dominating set is NP-complete for general graphs, bipartite graphs [23], and split graphs [16], whereas it is solvable in polynomial time for trees [17].
@@
The domatic number of a graph was introduced by Cockayne and Hedetniemi [4]. Furthermore, many variants of this concept were introduced and studied by B. Zelinka [31]. One such variation is the total domatic number, which was introduced by Cockayne et al. [3]. Several works have been done on the total domatic number [29, 30]. In 2024, Paul and Santra introduced three variations of transitivity, namely, $2$-transitivity [21], $d_{2}$-transitivity [20], and strong transitivity [22]. The $2$-transitivity, $d_{2}$-transitivity, and strong transitivity are based on $2$-domination, $d_{2}$-domination, and strong domination, respectively. In this article, we explore another natural variation of the transitivity problem, focusing on total domination instead of domination. We introduce such a variation of the transitivity problem, namely total transitive partition. Total transitivity $Tr_{t}(G)$ equals the maximum order of a vertex partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$ of $G$ obtained by repeatedly removing a total dominating set from $G$ until no vertices remain. Thus, $V_{1}$ is a total dominating set of $G$, $V_{2}$ is a total dominating set of the graph $G_{1}=G-V_{1}$, and, for $12\leq i\leq k-$ 1, $V_{i+1}$ is a total dominating set in the graph $G_{i}=G-\bigcup_{j=1}^{i}V_{i}$. A vertex partition of order $Tr_{t}(G)$ is called a $Tr_{t}$-partition. Note that every total transitive partition is also a transitive partition. Therefore, for any graph $G$, $1\leq Tr_{t}(G)\leq Tr(G)\leq n$. Since the total dominating set exists only when $G$ does not have any isolated vertices, in this paper, we focus only on graphs with no isolated vertices.
@@
An equivalent definition of total transitivity can be given in terms of total domination among partite sets and the set itself. A total transitive partition of size $k$ is defined as a partition of the vertex set into $k$ parts, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, such that for all $1\leq i\leq j\leq k$, $V_{i}$ totally dominates $V_{j}$. The maximum order of such a total transitive partition is called the total transitivity of $G$ and is denoted by $Tr_{t}(G)$.
@@
We define another variation of transitivity, namely weak total transitivity. A weak total transitive partition of size $k$ is defined as a partition of the vertex set into $k$ parts, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, such that for all $1\leq i\leq j\leq k$ and $i\neq k$, $V_{i}$ totally dominates $V_{j}$. The maximum order of such a weak total transitive partition is called the weak total transitivity of $G$ and is denoted by $Tr_{t}^{w}(G)$. Later, we will use this parameter to solve total transitivity in a tree. Clearly, if $\pi$ is a total transitive partition, it is also a weak total transitive partition, and hence, for a graph $G$, $Tr_{t}(G)\leq Tr_{t}^{w}(G)$. Also, note that in a weak total transitive partition $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, $V_{k}$ may not be a total dominating set of $G-\bigcup_{j=1}^{k-1}V_{i}$ but if $\pi$ is a total transitive partition, $V_{k}$ is a total dominating set of $G-\bigcup_{j=1}^{k-1}V_{i}$.
@@
    Given a split graph $G$ and $11\leq p\leq\omega(G)-$ 1, we provide necessary conditions for $Tr_{t}(G)=p$.
@@
    We prove that Maximum Total Transitivity Problem can be solved in linear time for bipartite chain graphs.
  5. (5.)
    We design a polynomial time algorithm for Maximum Total Transitivity Problem in trees.
The rest of the paper is organized as follows. Section 2 contains basic definitions and notations that are followed throughout the article. This section also discusses the properties of total transitivity of graphs. In section 3, we characterize split graphs with total transitivity equal to $1$ and $\omega(G)-1$. Also, in this section, we provide necessary conditions for $Tr_{t}(G)=p$ for split graphs $G$ and $11\leq p\leq\omega(G)-$ 1. Section 4 shows that the Maximum Total Transitivity Decision Problem is NP-complete in bipartite graphs. In section 5 we prove that Maximum Total Transitivity Problem can be solved in linear time for bipartite chain graphs. In Section 6, we design a polynomial-time algorithm for solving Maximum Total Transitivity Problem in trees. Finally, Section 7 concludes the article with some open problems.
@@ 2.1 Notation and definition
Let $G=(V,E)$ be a graph with $V$ and $E$ as its vertex and edge sets, respectively. A graph $H=(V^{\prime},E^{\prime})$ is said to be a subgraph of a graph $G=(V,E)$ if and only if $V^{\prime}\subseteq V$ and $E^{\prime}\subseteq E$. For a subset $S\subseteq V$, the induced subgraph on $S$ of $G$ is defined as the subgraph of $G$ whose vertex set is $S$ and whose edge set consists of all of the edges in $E$ that have both endpoints in $S$, and it is denoted by $G[S]$. The complement of a graph $G=(V,E)$ is the graph $\overline{G}=(\overline{V},\overline{E})$, such that $\overline{V}=V$ and $\overline{E}=\{uv|uv\notin E\text{ and }u\neq v\}$. The open neighbourhood of a vertex $x\in V$ is the set of vertices $y$ adjacent to $x$, denoted by $N_{G}(x)$. The closed neighborhood of a vertex $x\in V$, denoted as $N_{G}[x]$, is defined by $N_{G}[x]=N_{G}(x)\cup\{x\}$. A subset $D\subseteq V$ is said to be a total dominating set of $G$, if for every vertex $x\in V$, there exists a vertex $y\in D$ such that $xy\in E$. Breadth first search (BFS) is a graph searching technique that systematically explores all the vertices of a connected graph. The process starts at a designated vertex and explores all its neighbours before moving on to the next level of neighbours. The order in which the vertices are visited is called the BFS ordering. A vertex is said to be a leaf vertex if its degree is $1$, and the vertex that is adjacent to a leaf vertex is said to be the support vertex of the leaf vertex.
@@
A bipartite graph $G=(X\cup Y,E)$ is called a bipartite chain graph if there exists an ordering of vertices of $X$ and $Y$, say $\sigma_{X}=(x_{1},x_{2},\ldots,x_{n_{1}})$ and $\sigma_{Y}=(y_{1},y_{2},\ldots,y_{n_{2}})$, such that $N(x_{n_{1}})\subseteq N(x_{n_{1}-1})\subseteq\ldots\subseteq N(x_{2})\subseteq N(x_{1})$ and $N(y_{n_{2}})\subseteq N(y_{n_{2}-1})\subseteq\ldots\subseteq N(y_{2})\subseteq N(y_{1})$. Such ordering of $X$ and $Y$ is called a chain ordering, and it can be computed in linear time [15].